# Frequency-Time Diffusion with Neural Cellular Automata

### v1

@@ Abstract
Denoising Diffusion Models (DDMs) have become the leading generative technique for synthesizing high-quality images but are often constrained by their UNet-based architectures that impose certain limitations. In particular, the considerable size of often tens to hundreds of millions of parameters makes them impractical when hardware resources are limited. However, even with powerful hardware, processing images in the gigapixel range is difficult. This is especially true in fields such as microscopy or satellite imaging, where such challenges arise from the limitation to a predefined generative size and the inefficient scaling to larger images. We present two variations of Neural Cellular Automata (NCA)-based DDM methods to address these challenges and jumpstart NCA-based DDMs: Diff-NCA and FourierDiff-NCA. Diff-NCA performs diffusion by using only local features of the underlying distribution, making it suitable for applications where local features are critical. To communicate global knowledge in image space, naive NCA setups require timesteps that increase with the image scale. This is problematic when global features are essential, such as when synthesizing CelebA-like images. We solve this bottleneck of current NCA architectures by introducing FourierDiff-NCA, which advances Diff-NCA by adding a Fourier-based diffusion process and combines the frequency-organized Fourier space with the image space. By initiating diffusion in the Fourier domain and finalizing it in the image space, FourierDiff-NCA accelerates global communication. We validate our techniques by using Diff-NCA (208k parameters) to generate high-resolution digital pathology scans at $\mathbf{576\times 576}$ resolution and FourierDiff-NCA (887k parameters) to synthesize CelebA images at $64\times 64$, outperforming VNCA and five times bigger UNet-based DDMs. In addition, we demonstrate FourierDiff-NCA’s capabilities in super-resolution, OOD image synthesis, and inpainting without additional training.
@@
Figure 1: Diff-NCA is parameter efficient while being able to generate infinite seamless images. The one-cell model size allows FourierDiff-NCA to be applied to different input sizes, thus generating images of different shapes and scales. This same architecture also allows it to efficiently regenerate parts of an image in an inpainting task and perform superresolution on an existing image.
Introduction
@@
Denoising Diffusion Models (DDMs) have emerged as the leading architecture for generating high-quality images (Dhariwal and Nichol 2021). However, UNet-based (Ronneberger, Fischer, and Brox 2015) DDMs have significant limitations, such as a model size of tens to hundreds of millions of parameters, that restrict their use in environments with limited hardware. But even when powerful hardware is available, the inability to adapt to different image sizes and the inefficiency when increasing the generative size makes synthesizing gigapixel images, such as digital pathology scans or satellite images, challenging.
@@
Neural Cellular Automata (NCAs) (Gilpin 2019) offer a promising architectural alternative to the UNet. Unlike traditional deep learning models, NCAs adopt a unique single-cell model architecture inspired by biological cell communication. They interact only with their immediate neighbors, keeping the model size small while efficiently encoding information. Local communication makes NCAs highly adaptable to different image sizes (Mordvintsev, Niklasson, and Randazzo 2021; Pajouheshgar et al. 2023). However, initial efforts at NCA-based image generation encountered limitations. Synthesizing beyond image sizes of $64\times 64$ faced obstacles due to the inherent need for local communication, which requires increasing steps for global knowledge communication. Increasing the step count not only slows down the run-time but also complicates the learning process and increases the required VRAM during training time (Kalkhof, González, and Mukhopadhyay 2023).
@@
We introduce two new DDM methods based on Neural Cellular Automata to address these challenges and jumpstart NCA-based DDMs: Diff-NCA and FourierDiff-NCA. Diff-NCA focuses on local features of the underlying distribution, making it suitable for applications where local details are crucial, such as digital pathology scans or satellite imagery. With merely 208k parameters, it can synthesize digital pathology scans, multiple magnitudes larger than the training size of $64\times 64$, which we demonstrate by synthetizing images of size $576\times 576$. We introduce FourierDiff-NCA to recognize the importance of global features in datasets such as CelebA (Liu et al. 2015). It uses a Fourier-based diffusion approach to connect the frequency-structured Fourier space to the image space by starting diffusion in the Fourier domain and completing it in the image space. This method solves global communication and simplifies the optimization path using a quarter of the Fourier space. FourierDiff-NCA can generate $64\times 64$ CelebA images with merely 887k parameters while maintaining global knowledge.
@@
In our evaluation, we compare our proposed FourierDiff-NCA with UNet-based DDMs and VNCA (Palm et al. 2022) on the CelebA dataset. FourierDiff-NCA achieves an FID score of 77.84 with merely 887k parameters, outperforming five times bigger UNets with 3.94m parameters and an FID score of 128.2 and the ten times bigger VNCA with an FID score of 299.9. Moreover, we demonstrate our architecture’s flexibility by highlighting its ability to perform super-resolution, inpainting, and synthesize images of out-of-distribution sizes, further supporting the potential of our proposed models for various image-generation tasks. Diff-NCA and FourierDiff-NCA represent a considerable advance in NCA-based image generation, jumpstarting NCA-based DDMs. They offer scalable, efficient, and adaptable solutions (illustrated in Figure 1) that overcome the limitations of traditional DDMs and provide a path to high-quality image synthesis at arbitrary scales on minimal hardware.
@@
Related Work
While Neural Cellular Automata can generalize across different image sizes, current NCA image generation approaches, such as VNCA (Palm et al. 2022) and GANCA (Otte et al. 2021), have not exploited this property. These methods are limited to $64\times 64$ pixel images and show constraints in their random sampling results. This unexplored potential motivates our work with Diff-NCA and FourierDiff-NCA.
Neural Cellular Automata (NCA)
NCAs differ substantially from conventional deep learning architectures as they are built on a one-cell model (a detailed introduction to NCAs can be found in the supplementary). This one-cell model is replicated in each cell of an image, where it communicates only with its direct neighbors. To gain global knowledge, this learned update rule is repeated multiple times, which means that a naive NCA device requires at least 100 steps to communicate across a $100\times 100$ image. Due to the size of the single-cell model, NCAs typically require below a million parameters to handle complex tasks such as growing an image from a single cell (Mordvintsev et al. 2020) or image generation (Otte et al. 2021; Palm et al. 2022). Although the model sizes are small, the VRAM requirements for NCAs increase exponentially during training as backpropagation requires all duplications of the NCA in the world and across all timesteps. This makes it challenging to train NCAs on a large number of steps (Kalkhof, González, and Mukhopadhyay 2023).
NCA Image Generation
Current NCA image generation methods are based on two main architectures: variational autoencoders (VAEs) (Kingma and Welling 2013) and generative adversarial networks (GANs) (Goodfellow et al. 2014).
GAN-based: GANCA (Otte et al. 2021) explores the generation of emojis with a size of $30\times 30$ pixels, using an NCA as a generator in a traditional GAN framework augmented by a classical discriminator.
VAE-based: VNCA (Palm et al. 2022) expands images up to $64\times 64$ pixels by integrating an NCA into the image generation component of a VAE (Kingma and Welling 2013). By distributing the latent vector of the VAE to each cell of the NCA, the method can regenerate the encoded image.
@@
Denoising Diffusion Models
Denoising Diffusion Models (Ho, Jain, and Abbeel 2020) are based on the idea of sequentially diffusing an unknown distribution into a univariate Gaussian (Sohl-Dickstein et al. 2015) distribution. Then, using adequately small diffusion steps, one can train a denoising model to inverse these steps. Eventually, new samples from the unknown target distribution can be generated by sampling from the known distribution and repeatedly applying the denoising model, sequentially pushing the sample back into the target distribution. Several incremental improvements have been proposed over the recent years (Song, Meng, and Ermon 2020; Nichol and Dhariwal 2021; Rombach et al. 2022), and this family of models has shown superior image generation quality compared to the GAN-based standard to date (Dhariwal and Nichol 2021).
Typically, the denoising model is a modified version of a UNet (Ronneberger, Fischer, and Brox 2015), which comes with several limitations: UNets are typically compute-heavy architectures with several millions of parameters. Further, their performance is bound to the training modalities and can not out-of-the-box produce images of varying scales or be used for tasks such as inpainting or superresolution. For such tasks, adapted architectures are required for training UNet-based DDMs (Rombach et al. 2022; Ho et al. 2022; Saharia et al. 2022; Li et al. 2022; Lugmayr et al. 2022). On the contrary, our proposed architecture is flexible for different input modalities, such as different scales, and it can be used out-of-the-box for tasks like inpainting and superresolution.
Methodology
@@
Fourier Space: Single-Step Global Communication
@@
When using Neural Cellular Automata, one of the major limitations is the number of steps required for the model to acquire global knowledge. Since NCAs communicate with their direct neighbors, communication across a $100\times 100$ image in a naive setup thus requires 100 steps. The inherent structure of the Fourier representation, where lower frequency data lives in the middle of a two-dimensional Fourier space, allows NCAs operating in this space to communicate globally across the entire image space in a single step. Additional iterations can be used to refine the transmitted signal. The possibility of instantaneous communication over the entire image space arises from applying an inverse Fast Fourier Transform following the initial NCA communication, whereby all relevant data in the limited Fourier space is transferred to the global scale.
@@
Only a fraction of the Fourier space is required for this process since a significant portion of the information in the Fourier space is assumed to be insignificant from a quality point of view (we further illustrate this in the supplementary material).
FourierDiff-NCA Architecture
We approach the challenges of globally coherent diffusion with NCAs with FourierDiff-NCA by incorporating image and Fourier space. We use the separate NCA models $m_{1}$ and $m_{2}$ for the image and Fourier space. The denoising process is illustrated in Figure 2.
@@
Diff-NCA: As a subset of the FourierDiff-NCA architecture, Diff-NCA uses only local communication. The denoising task $*n*$ follows the procedure of DDMs and serves as the input of Diff-NCA. It is a combination of the input image $i$ and the noise $n$. An embedding $e$ includes the position $x,y$ of the NCA, the timestep $t$ of diffusion, and the NCA timestep. Diff-NCA predicts the noise $\bar{n}$ from $*n*$ by iterating $m_{1}$, $s$ times over the image, incrementing its perceptive range by 1 per step. The loss is computed between $\bar{n}$ and $n$ using a combined L2 and L1 loss, which diverges from the original L2 loss, to enhance convergence.
Embedding: The embedding’s linear information of pos x, y, diffusion timestep, and NCA timestep is processed using sinusoidal encoding (Vaswani et al. 2017), as introduced by DDM (Ho, Jain, and Abbeel 2020). A sequence of a linear layer of size 256, SiLU activation, and another linear layer mapping to 4 output channels is utilized. Our setup provides this information per NCA cell, thus using a two-dimensional encoding.
@@
FourierDiff-NCA: extends Diff-NCA by initiating the diffusion process by gathering global information in Fourier space. Through a Fast Fourier Transform (FFT) on $*n*$, FourierDiff-NCA receives the diffusion task $f$ in the Fourier space. We extract a $16\times 16$ cell window starting at the center of $f$ paired with the embedding $e$, simplifying communication. We use a $16\times 16$ quarter of the Fourier space, which contains enough details for global communication. After 32 iterations (required to communicate once across and back of a $16\times 16$ world) of $m_{2}$ in Fourier space, an inverse FFT is performed to convert back to the image space, and the process transitions to Diff-NCA, carrying all the information in the channels $c$ of the NCA.
Optimized Inference: When sampling with FourierDiff-NCA, shortcomings are observed in low- and medium-frequency information due to the limited ability of NCA to incorporate an independent update rule dependent on the diffusion step. This can lead to deviations from the planned diffusion schedule during inference. We propose an optimized inference method where the current step is predicted based on the average statistical properties of the current noisy image. This is compared to the expected denoising path of the training images. The statistical prediction is performed by computing the standard deviation of the first derivatives in the x and y directions. Given the increasing ambiguity of the step towards the end of the training process, we stick to the prescribed path when the maximum divergence of the last step (in the training set) exceeds the minimum noise of an earlier step.
Model Architecture
@@
We design our model, illustrated in Figure 2, with simplicity in mind. We keep the architecture identical for the image and Fourier space. Although, there is a difference in the number of input channels $c$, since converting data to the Fourier space results in two values per channel.
@@
Local communication is implemented through a $3\times 3$ 2D convolution. The output of that convolution is concatenation with the previous internal state $v$. The next layer is a $1\times 1$ 2D convolution that maps the concatenated vector to a hidden vector of depth $h=512$. Group normalization and a leaky ReLU activation are applied to ensure normalization and introduce nonlinearity.
Another $1\times 1$ 2D convolution then maps the hidden layer back to the channel size $c$, resulting in an output vector $o$. The internal state $v$ is updated by: $v\leftarrow v+o$, controlled by a random reset mechanism referred to as the fire rate, which is set to a probability of 50%. When the reset mechanism is activated, the update of the cells in question is set to 0.
Experimental Results

### v2

@@ Abstract
Despite considerable success, large Denoising Diffusion Models (DDMs) with UNet backbone pose practical challenges, particularly on limited hardware and in processing gigapixel images. To address these limitations, we introduce two Neural Cellular Automata (NCA)-based DDMs: Diff-NCA and FourierDiff-NCA. Capitalizing on the local communication capabilities of NCA, Diff-NCA significantly reduces the parameter counts of NCA-based DDMs. Integrating Fourier-based diffusion enables global communication early in the diffusion process. This feature is particularly valuable in synthesizing complex images with important global features, such as the CelebA dataset. We demonstrate that even a 331k parameter Diff-NCA can generate $512\times 512$ pathology slices, while FourierDiff-NCA (1.1m parameters) reaches a three times lower FID score of 43.86, compared to the four times bigger UNet (3.94m parameters) with a score of 128.2. Additionally, FourierDiff-NCA can perform diverse tasks such as super-resolution, out-of-distribution image synthesis, and inpainting without explicit training.
Machine Learning, ICML
@@
Figure 1: Diff-NCA is parameter efficient while being able to generate infinite seamless images. The one-cell model size allows FourierDiff-NCA to be applied to inputs, different from the training size, thus generating images of different shapes and scales. This same architecture also allows it to efficiently regenerate parts of an image in an inpainting task and perform superresolution on an existing image, without the need for retraining.
1 Introduction
@@
Denoising Diffusion Models (DDMs) have emerged as the leading architecture for generating high-quality images (Dhariwal & Nichol, 2021). However, UNet-based (Ronneberger et al., 2015) DDMs have significant limitations, such as a model size of tens to hundreds of millions of parameters, that restrict their use in environments with limited hardware. But even when powerful hardware is available, the inability to adapt to different image sizes and the inefficiency when increasing the generative size makes synthesizing gigapixel images, such as digital pathology scans or satellite images, challenging.
@@
Neural Cellular Automata (NCAs) (Mordvintsev et al., 2020; Gilpin, 2019) offer a promising architectural alternative to the UNet. Unlike traditional deep learning models, NCAs adopt a unique single-cell model architecture inspired by biological cell communication. They interact only with their immediate neighbors, keeping the model size small while efficiently encoding information. Local communication makes NCAs highly adaptable to different image sizes (Mordvintsev et al., 2021; Pajouheshgar et al., 2023). However, initial efforts at NCA-based image generation encountered limitations. Synthesizing beyond image sizes of $64\times 64$ faced obstacles due to the inherent need for local communication, which requires increasing steps for global knowledge communication. Increasing the step count not only slows down the run-time but also complicates the learning process and increases the required VRAM during training time (Kalkhof et al., 2023).
@@
We introduce two new DDM methods based on Neural Cellular Automata to address these challenges and jumpstart NCA-based DDMs: Diff-NCA and FourierDiff-NCA. Diff-NCA focuses on local features of the underlying distribution, making it suitable for applications where local details are crucial, such as digital pathology scans or satellite imagery. With merely 331k parameters, it can synthesize digital pathology scans, multiple magnitudes larger than the training size of $64\times 64$, which we demonstrate by synthetizing images of size $512\times 512$. We introduce FourierDiff-NCA to recognize the importance of global features in datasets such as CelebA (Liu et al., 2015). It uses a Fourier-based diffusion approach to connect the frequency-structured Fourier space to the image space by starting diffusion in the Fourier domain and completing it in the image space. This method solves global communication and simplifies the optimization path using a quarter of the Fourier space. FourierDiff-NCA can generate $64\times 64$ CelebA images with merely 1.1m parameters while maintaining global knowledge.
@@
In our evaluation, we compare our proposed FourierDiff-NCA with UNet-based DDMs and VNCA (Palm et al., 2022) on the CelebA dataset. FourierDiff-NCA achieves an FID score of 43.86 with merely 1.1m parameters, outperforming four times bigger UNets with 3.94m parameters and an FID score of 128.2 and the ten times bigger VNCA with an FID score of 299.9. Moreover, we demonstrate our architecture’s flexibility by highlighting its ability to perform super-resolution, inpainting, and synthesize images of out-of-distribution sizes, further supporting the potential of our proposed models for various image-generation tasks. Diff-NCA and FourierDiff-NCA represent a considerable advance in NCA-based image generation, jumpstarting NCA-based DDMs. They offer scalable, efficient, and adaptable solutions (illustrated in Figure 1) that overcome the limitations of traditional DDMs and provide a path to high-quality image synthesis at arbitrary scales on minimal hardware.
@@
2 Related Work
While Neural Cellular Automata can generalize across different image sizes, current NCA image generation approaches, such as VNCA (Palm et al., 2022) and GANCA (Otte et al., 2021), have not exploited this property. These methods are limited to $64\times 64$ pixel images and show constraints in their random sampling results. This unexplored potential motivates our work with Diff-NCA and FourierDiff-NCA.
2.1 Neural Cellular Automata (NCA)
NCAs differ substantially from conventional deep learning architectures as they are built on a one-cell model (a detailed introduction to NCAs can be found in the appendix A). This one-cell model is replicated in each cell of an image, where it communicates only with its direct neighbors. To gain global knowledge, this learned update rule is repeated multiple times, which means that a naive NCA requires at least 100 steps to communicate across a $100\times 100$ image. Due to the size of the single-cell model, NCAs typically require below a million parameters to handle complex tasks such as growing an image from a single cell (Mordvintsev et al., 2020) or image generation (Otte et al., 2021; Palm et al., 2022). Although the model sizes are small, the VRAM requirements for NCAs increase exponentially during training as backpropagation requires all duplications of the NCA in the world and across all timesteps. This makes it challenging to train NCAs on a large number of steps (Kalkhof et al., 2023).
2.2 NCA Image Generation
Current NCA image generation methods are based on two main architectures: variational autoencoders (VAEs) (Kingma & Welling, 2013) and generative adversarial networks (GANs) (Goodfellow et al., 2014).
GAN-based: GANCA (Otte et al., 2021) explores the generation of emojis with a size of $30\times 30$ pixels, using an NCA as a generator in a traditional GAN framework augmented by a classical discriminator.
VAE-based: VNCA (Palm et al., 2022) expands images up to $64\times 64$ pixels by integrating an NCA into the image generation component of a VAE (Kingma & Welling, 2013). By distributing the latent vector of the VAE to each cell of the NCA, the method can regenerate the encoded image.
@@ 2.2 NCA Image Generation
2.3 Denoising Diffusion Models
Denoising Diffusion Models (Ho et al., 2020) are based on the idea of sequentially diffusing an unknown distribution into a univariate Gaussian (Sohl-Dickstein et al., 2015) distribution. Then, using adequately small diffusion steps, one can train a denoising model to inverse these steps. Eventually, new samples from the unknown target distribution can be generated by sampling from the known distribution and repeatedly applying the denoising model, sequentially pushing the sample back into the target distribution. Several incremental improvements have been proposed over the recent years (Song et al., 2020; Nichol & Dhariwal, 2021; Rombach et al., 2022; Hoogeboom et al., 2023; Croitoru et al., 2023; Gao et al., 2023), and this family of models has shown superior image generation quality compared to the GAN-based standard to date (Dhariwal & Nichol, 2021).
Typically, the denoising model is a modified version of a UNet (Ronneberger et al., 2015), which comes with several limitations: UNets are typically compute-heavy architectures with several millions of parameters. Further, their performance is bound to the training modalities and can not out-of-the-box produce images of varying scales or be used for tasks such as inpainting or superresolution. For such tasks, adapted architectures are required for training UNet-based DDMs (Rombach et al., 2022; Ho et al., 2022; Saharia et al., 2022; Li et al., 2022; Lugmayr et al., 2022). On the contrary, our proposed architecture is flexible for different input modalities, such as different scales, and it can be used out-of-the-box for tasks like inpainting and superresolution.
3 Methodology
@@
3.1 Fourier Space: Single-Step Global Communication
@@
When using Neural Cellular Automata, one of the major limitations is the number of steps required for the model to acquire global knowledge. Since NCAs communicate with their direct neighbors, communication across a $100\times 100$ image in a naive setup thus requires 100 steps. The inherent structure of the Fourier representation (we explain this in-depth in the appendix B), where lower frequency data lives in the middle of a two-dimensional Fourier space, enables a fundamental shift in the communication pattern. This shift allows NCAs operating in this space to achieve global communication across the entire image in a single step, a stark contrast to the linear step-wise progression needed in image space. Additional iterations can be used to refine the transmitted signal. The possibility of instantaneous communication over the entire image space is a significant advantage, that arises from applying an inverse Fast Fourier Transform following the initial NCA communication, whereby all relevant data in the limited Fourier space is transferred to the global scale. After the initial acquisition of global information in Fourier space, each cell in the image space starts from a comprehensive understanding of the global context and then adjusts the details based on local information, which is a clear departure from the traditional ”detail-first” approach.
@@
Notably, only a fraction of the Fourier space is required for this process since a significant portion of the information in the Fourier space is assumed to be insignificant from a quality point of view (we further illustrate this in the appendix in Figure 11).
3.2 FourierDiff-NCA Architecture
In FourierDiff-NCA, we address the challenge of achieving global coherence by beginning the diffusion process in Fourier space to capture global information and then transitioning to image space to integrate local details. This approach allows us to combine global and local information effectively without the need for a high number of NCA steps. We use the separate NCA models $m_{1}$ and $m_{2}$ for the image and Fourier space. The denoising process is illustrated in Figure 2.
@@ 3.2 FourierDiff-NCA Architecture
Diff-NCA: As a subset of the FourierDiff-NCA architecture, Diff-NCA uses only local communication (illustrated in the appendix in Figure 10). It is noteworthy that Diff-NCA can be run independently of FourierDiff-NCA, resulting in a diffusion process based solely on the image space and thus considering only local features. The denoising task $i_{n}$ follows the procedure of DDMs and serves as the input of Diff-NCA. It is a combination of the input image $i$ and the noise $n$. An embedding includes the position $x,y$ of the NCA, the timestep $t$ of diffusion, and the NCA timestep. Diff-NCA predicts the noise $n_{p}$ from $i_{n}$ by iterating $m_{1}$, $s$ times over the image, incrementing its perceptive range by 1 per step. The loss is computed between $n_{p}$ and $n$ using a combined L2 and L1 loss, which diverges from the original L2 loss, to enhance convergence.
[…]
@@
FourierDiff-NCA: extends Diff-NCA by initiating the diffusion process by gathering global information in Fourier space. Through a Fast Fourier Transform (FFT) on $i_{n}$, FourierDiff-NCA receives the diffusion task $f$ in the Fourier space. We extract a $16\times 16$ cell window starting at the center of $f$ paired with the embedding $e$, simplifying communication. This $16\times 16$ quarter of the Fourier space, contains enough details for global communication. After 32 iterations (required to communicate once across and back of a $16\times 16$ world) of $m_{2}$ in Fourier space, an inverse FFT is performed to convert back to the image space, and the process transitions to Diff-NCA, providing initial global information in the channels $c$ of each NCA cell.
3.3 Model Architecture
@@
We design our model, illustrated in Figure 2, with simplicity in mind. We keep the architecture identical for the image and Fourier space. However, there is a difference in the number of input channels $c$, since converting data to the Fourier space results in two values per channel.
@@
Local communication is implemented through a $3\times 3$ 2D convolution. The output of that convolution is concatenation with the previous internal state $v$ and the embedding $e$. This concatenated vector is multiplied with the output of $M_{eb1}$. The next layer is a $1\times 1$ 2D convolution that maps the concatenated vector to a hidden vector of depth $h=512$. Group normalization and a leaky ReLU activation are applied to ensure normalization and introduce nonlinearity.
Now the output of the ReLU is multiplied by the output of $M_{eb2}$. Afterwards, another $1\times 1$ 2D convolution maps the hidden layer back to the channel size $c$, resulting in an output vector $o$. The internal state $v$ is updated by: $v\leftarrow v+o$, controlled by a random reset mechanism referred to as the fire rate, which is set to a probability of 90%. When the reset mechanism is activated, the update of the cells in question is set to 0.
Following practices aligned with the leading methods in the field (Dhariwal & Nichol, 2021; Ho et al., 2022; Meng et al., 2023), our model incorporates an exponential moving average (EMA) on the weights, with a decay rate of 0.99.
Figure 3: Qualitative comparison between FourierDiff-NCA (1.85m), VNCA, and DDM based on UNet M. With a parameter count of 1.85m, 9.73m, and 3.94m, respectively.
Figure 4: Influence of parameter count to image generation performance for FourierDiff-NCA, UNet, and VNCA (the detailed numbers can be found in the appendix in Table 2). Green quadrant: low parameters, high performance; red quadrant: high parameters, low performance; yellow quadrants: tradeoff.
4 Experimental Results